# Mind Your Format: Towards Consistent Evaluation of In-Context Learning Improvements

### v1

@@ 5.3 Transfer Between Prediction Methods
As discussed in Section 4.2, no prediction method consitently outperforms others across all models and datasets. Therefore, to find an optimal setup for a new ICL improvement, one needs to evaluate every prediction technique in multiple templates. We investigate whether it is possible to find the universally optimal prompt for different methods to reduce the total computational cost.
@@
Table 4: Intersection-over-Union between different decoding methods averaged over 19 models results obtained in Random 2-shot setup.
To answer this question, we calculate the IoU between top 10 performing templates for each method for a fixed set of demonstrations. Results in Table 4, averaged over 19 models display that transfer between prediction methods is substantially low. Consequently, the prompt format sensitivity issue incurs on practitioners the burden of tuning templates for every prediction method they may wish to employ in conjunction with their own approach.
5.4 Transfer Between Example Selection Methods
Having found that the best performing templates are specific both to a model and a prediction method, we set out to find whether the best formats are transferable at least between demonstration sets in the same setup. Similarly to previous experiments, we calculate IoU for 10 templates yielded the highest scores for every method.
Figure 5: IoU of 10 best templates for each example selection method on AG News dataset. Method-N indicates that Method was used to select $N$ examples.
Results in Figure 5 illustrate that simply adding several demonstrations, even if they were obtained with the same method, highly disturbs the ranking of the best templates. This justifies the necessity to evaluate example selection methods on a range of templates to avoid misinterpretation of the results.
Based on the above, we conclude that the results of evaluation of various ICL improvements without consideration of template sensitivity issue are hardly reliable for several reasons. First, as the best templates do not transfer between models even within the same model family, scoring a method across several models using the same format will inevitably lead to underestimation of the method for all models except the one for which the format was tuned. Next, as there is no transferability between settings, the format selection procedure has to be precisely described and applied for all evaluated settings for a fair comparison. Moreover, a comparison is unfair even for a non-optimized template, as it introduces an element of chance: a randomly selected template might perform well for one setup but poorly for another.
To sum up, there are no universally good performing prompt formats. Therefore, the results of different methods evaluation can be reliable only if they are aggregated over several templates or if the best performing template was selected for every setting in the same manner. The former approach requires accounting for variance of the scores and makes comparison less apparent, while the latter is remarkably computationally expensive.
@@ 6 Template Ensembles
Results of Section 5 indicate that finding an optimal template for a given model is insufficient as it cannot be transferred between different prediction methods, demonstration selections, their order and number. At the same time, tuning prompt template for a given setting is extremely computationally expensive. As a first step towards mitigating this issue we propose to ensemble model predictions over several templates. This approach is widely used in machine learning Ho (1995); Lakshminarayanan et al. (2017) for improving generalization and increasing average performance.
Formally, the algorithm computes label probabilities across predictions for each of $N$ templates, where $N$ is the ensemble size, and outputs the label with the highest average probability.
Model Direct Channel Calibrate
@@
We begin with determining the minimal ensemble size that consistently reduces variance while keeping average performance on the same level or higher. We observe that for the majority of models and prediction methods an ensemble achieves the best accuracy when its size reaches 4 or 5 (see example in Figure 6), with further expansion being less effective. We also find empirical evidence that smaller ensembles may demonstrate unstable behavior, with the possibility of a drop in performance if a suboptimal template is randomly sampled. For the sake of uniformity, in  Table 5 we report results for ensembles of size 5.
Next, we evaluate performance gains introduced by Template Ensemble for different prediction methods. Our findings in Table 5 and Appendix J indicate that ensembles increase the resulting accuracy for all evaluated models and prediction methods. Most importantly, variance is significantly reduced for most models and methods. Therefore, we can conclude that our method allows to preserve an increase in average performance provided by various ICL advancements methods while effectively mitigating the template sensitivity issue.
Figure 6: Ensemble size ablation for Falcon 40B 2-shot learning performance on the SST-2 dataset.
@@ 7 Conclusion
While we propose Template Ensembles as an initial solution to this problem, the sensitivity of language models to minor prompt variations is yet to be addressed. Consequently, we believe that the research community should take this problem into account when developing new models, evaluation benchmarks, or in-context learning methods.
@@
Due to limited computational resources and the high cost for evaluation on a large range of models, we only focus on four classification datasets. Moreover, we only compare two example selection methods to a random selection baseline, potentially overlooking other effective approaches.
Additionally, the space of templates could be expanded for more comprehensive experimentation. For example, we did not explore label-mapping, including random labelling, which is an important aspect of the template.
@@
Finally, we only evaluate decoder-only models pre-trained on next token prediction objective. We leave other architectures as well as instruction-fine-tuned models to the future work.
@@ Appendix A Templates from Prior Work
Tables 6, 7, 8 and 9 provide a comparison of all the templates used in the works presenting all methods we evaluate. Noticeably, prompt formats (and the choice of label words for some formats and datasets) used in works proposing investigated methods have no intersection at all. This is additionally concerning, since the original papers proposing these methods cite each other, signifying that the use of different template pools can be intentional. For instance, Channel method outperforming Calibration in the evaluation conducted in Min et al. (2022b) might be explained by selecting a more favorable set of templates for the method proposed in the paper rather than by the superiority of the method itself.
@@
Appendix B Models Selection
Our initial pool of models for investigation consisted of 23 models. We evaluated each of them in 0- and 2-shot in three prediction methods investigated in our work on each of four datasets resulting in 12 runs. For each of 12 runs in both 0- and 2-shot we compare model’s performance averaged over 10 random templates.
Based on the results presented in Table 10, we restricted the final pool of models for evaluation by those that consistently outperform 0-shot in 2-shot setting, i.e. demonstrate performance boost from ICL. More specifically, we keep the models that had 8 or more wins in 2-shot against 0-shot.
Table 10: Evaluation of advanced prediction methods for all models on 4 datasets in 0- and Random-2-shot. Models that were removed from further evaluation are highlighted in gray.
@@ Appendix B Models Selection
Table 10: Evaluation of advanced prediction methods for all models on 4 datasets in 0- and Random-2-shot. Models that were removed from further evaluation are highlighted in gray.
@@ Appendix C Full Baseline Results
Table 11 presents the evaluation of all 19 models in the default setting in 0/2/4-shot on 4 datasets. These results illustrate that the template sensitivity issue is inherent in all models despite their size, and is not efficiently mitigated with the increase of number of demonstrations.
@@
Table 11: Classification accuracy of all models on all datasets in the default setting. The results are aggregated over 10 templates for each set of random demonstrations, i.e. 10 runs for 0-shot and 30 runs for few-shot.
@@ Appendix D Template Parts Analysis
(a) Direct
(b) Channel
(c) Calibration
Figure 7: Accuracy scores of template parts on the SST-2 dataset, with Random 2-shot for all prediction methods
In addition to transferability investigation, we analyze what impact each part of a template – input verbalizer, output verbalizer, intra- and inter-separators – brings on performance. For instance, it is possible that the inclusion of certain verbalizers in a template consistently leads to a decline in accuracy, irrespective of the other components.
To find that out, we decompose templates back into components and measure the distribution of scores for the different variations of each component separately. The results presented in  Figure 7 illustrate that even for SoTA models (LLaMA 2 70B and Falcon 40B) many verbalizers exhibit high variance. In other words, even if we discovered that certain template yields good performance for a fixed setup, it is neither guaranteed to transfer to other setups nor even ensure that changing single component will not have detrimental effects.
In addition to non-transferability of whole templates, we notice that individual verbalizers also exhibit non-transferability, both between models and prediction methods. For instance, while “It was {}“ ranks highest among output verbalizers for LLaMA 2 70B with Direct prediction method, it’s one of the worst for Falcon 40B.
@@
To summarize, we highlight the complex interactions between the various components of a template and their influence on model performance. We propose that both whole templates and individual verbalizers transferability is limited and require the further investigation.
@@ Appendix E Prediction Methods
We provide the results of advance prediction methods evaluation for all models for both 0- and Random-2-shot settings in Table 10. We conclude from this comparison that neither of the advanced prediction strategies do not curb prompt format sensitivity consistently across models and datasets. Moreover, when accounting for the spread in accuracy scores caused by this issue, the advantages of these methods over Direct become elusive.
Appendix F Example Selection Methods
Full results of evaluation of various demonstration selection techniques in 4-shot using Direct method are presented in Table 12.
@@ Appendix F Example Selection Methods
The evaluation highlights that intricate example selection techniques rarely significantly outperform the default random selection baseline when evaluated on a set of random templates. One would argue that the prompt format choice is inseparable from the method itself and thus such comparison is invalid. However, since the best performing formats cannot be transferred neither between different models, nor even between the sets of different sizes selected with the same method, a proper evaluation of each method would require finding the best template for each setup. Not only this procedure is computationally expensive, but sometimes is simply impossible as authors of proposed methods frequently omit describing their prompt selection algorithm.
Appendix G Accuracy Deterioration
We plot the dependence of accuracy on the rank of templates in Figure 8. The results are aggregated across 19 models. Each model was evaluated on 30 random templates with Direct prediction method and the same set of 2 randomly selected demonstrations. We observe that for SST-2 and AGNews datasets the mean quality of the 10th-best template is around 0.9 of the best template score, which we consider a successful transfer. Despite the more rapid decay for the DBPedia-14 and TREC, considering variation across models, we still count first 10 formats as performing on par with the best one.
Figure 8: Relative quality of templates sorted by their resulting classification accuracy.
Appendix H Transfer evaluation by Spearman rank correlation
One of the possible means to evaluate template transfer is to calculate Spearman rank correlation between scores of all templates. As can be seen from Figure 9 this method yields higher correlations than calculating IoU over 10 best formats, but transferrability is still far from perfect (e.g. for SST-2 and TREC datasets).
Figure 9: Spearman correlation rank of 30 templates for models evaluated in Direct-Random-2-shot setting.
Appendix I IoU Between Models
@@
Similarly to Figure 4 we provide Intersection-over-Union of 10 best prompt formats for all 19 models on all datasets explored in our work in Figure 10. These heatmaps illustrate that transfer of the best performing templates between models is extremely low for all datasets.
Appendix J Ensembles
Tables 13 and 14 show the results of Template Ensemble evaluation on a broader set of models and datasets. For most setups, ensemble of size 5 exhibited better performance than a single template.
@@
Table 13: Comparison of models’ 2-shot learning performance on the SST-2 dataset using 5-templates ensembles against a single template. Results are averaged over 5 random seeds.
@@
Table 14: Comparison of models’ 2-shot learning performance on the TREC dataset using 5-templates ensembles against a single template. Results are averaged over 5 random seeds.

### v3

@@ 5.3 Transfer Between Prediction Methods
As discussed in Section 4.2, no prediction method that we evaluate can consistently outperform others across all models and datasets. Therefore, to find an optimal setup for a new ICL improvement, one needs to evaluate every prediction technique in multiple templates. We investigate the possibility of finding a universally optimal prompt for different methods to reduce the total computational cost.
@@
Table 4: Intersection-over-Union for pairs of prompting methods averaged over the results of 19 base models obtained in the Random 2-shot setup. Standard deviations are in subscript.
To answer this question, we calculate the IoU between top 10 performing templates for each method for a fixed set of demonstrations. Results in Table 4 display that similarly to the models, the transfer between prediction methods is also low. Consequently, the prompt format sensitivity issue creates a burden on authors of new ICL modifications; they must tune templates for every prediction method they want to combine with their own approach.
5.4 Transfer Between Demonstration Selection Methods
Having found that the best-performing templates are specific both to the model and the prediction method, we now aim to find whether the best formats would be the same for different demonstration sets in the same setup. Similarly to previous experiments, we calculate IoU for 10 templates that yield the highest scores for each method.
Results in Figure 5 illustrate that simply adding demonstrations, even if they were obtained with the same method, can significantly alter the ranking of the best templates. This justifies the necessity to evaluate example selection methods on a range of templates to avoid misinterpretation of the results.
Figure 5: IoU of 10 best templates for example selection methods on the AG News dataset. Method-N indicates that Method was used to select $N$ examples.
5.5 Discussion
Based on the above findings, we conclude that the results of evaluation of various ICL improvements without consideration of template sensitivity issue are hardly reliable for several reasons. First, as the best templates do not transfer between models even within the same family, scoring a method across several models using the same format will inevitably lead to underestimation of the method for all models except the one for which the format was tuned. Next, as there is little evidence of transfer between setups, the format selection procedure needs to be precisely described and applied in all evaluated settings for a fair comparison.
In summary, we find that there are no universally well-performing prompt formats. Therefore, the results of in-context learning evaluation can be reliable only if they are aggregated over several templates or if each setting is evaluated in its best-performing template. The former approach requires accounting for the variance of the scores and makes comparison less apparent, while the latter can be computationally expensive.
@@ 6 Template Ensembles
Model Direct Channel Calibration
@@
Table 5: Comparison of 2-shot learning performance on the SST-2 dataset using ensembles of 5 templates and a single template. Standard deviations over 5 random seeds are in subscript, best accuracy for each model is in bold.
To reduce the variance in performance caused by the template choice, we propose to ensemble model predictions across multiple templates. This approach is widely used in machine learning Ho (1995); Lakshminarayanan et al. (2017) for improving the predictive performance of the model, as well as its robustness, and can be viewed as a form of test-time augmentation Krizhevsky et al. (2012); Simonyan and Zisserman (2015). Prior work on prompt ensembling has shown significant gains by training a boosting algorithm on model outputs Hou et al. (2023); by contrast, our method needs only the pretrained model predictions without additional training.
Formally, our method computes label probabilities across predictions for each of $N$ templates, where $N$ is the ensemble size, and outputs the label with the highest average probability. In early experiments, we tried selecting the most common label among the predictions; however, we found this voting strategy to perform poorly on tasks with a large number of classes. It is also important to note that ensembling $N$ predictions involves running the model $N$ times more compared to single-format evaluation, which makes this approach more computationally expensive. We view template ensembles as a baseline solution for the problem of prompt format sensitivity and leave the exploration of more efficient methods to future work.
Figure 6: Template ensemble accuracy as a function of its size for Falcon 40B on the SST-2 dataset in the 2-shot learning setup. Dashed lines depict the results of baseline methods averaged over 10 templates.
We begin with determining the minimal ensemble size that consistently reduces variance while increasing the average performance. We observe that for the majority of models and prediction methods, an ensemble achieves the best accuracy when its size reaches 4 or 5 (see an example in Figure 6), with further expansion being less effective. We also found that smaller ensembles may demonstrate unstable behavior, with the possibility of a drop in performance if a suboptimal template is sampled. Therefore, we report results for ensembles of size 5 and average the results over 5 random seeds.
Next, we evaluate the performance gains of Template Ensembles for different prediction methods. Our findings in Tables 5 and 20 and Appendices K and L indicate that ensembles increase the accuracy for all evaluated models and prediction methods. Most importantly, they also significantly reduce the variance caused by the template choice for most setups. Therefore, we conclude that template ensembling allows to preserve the increase in accuracy provided by ICL modifications while mitigating the template sensitivity issue.
@@ 7 Conclusion
While we propose Template Ensembles as an initial solution to this problem, the general sensitivity of language models to minor prompt variations is yet to be addressed. Consequently, we believe that the research community should take this problem into account when developing new models, evaluation benchmarks, or in-context learning methods.
@@
Due to limited computational resources and the high cost for evaluation on a large range of models, we only focus on four classification datasets. Moreover, we only compare two example selection methods to a random baseline, potentially overlooking other effective approaches.
Additionally, the space of templates could be expanded for more comprehensive experimentation. For example, we did not explore label mapping, including random labels, which is an important aspect of the template.
@@ Appendix A Templates from Prior Work
Tables 6, 7, 8 and 9 provide a comparison of all the templates used in the works presenting all methods we evaluate. Noticeably, prompt formats (and the choice of label words for some formats and datasets) used in works proposing investigated methods have no intersection. This is also concerning, since the original papers proposing these methods refer to each other. For instance, Channel prompting outperforming Calibration in Min et al. (2022b) might be explained by selecting a more favorable set of templates for the method proposed in the paper rather than by the advantages of the method itself.
@@
Appendix B Model Selection
Our initial evaluation pool consisted of 23 models. We evaluated each of them in 0-shot and 2-shot settings with three prediction methods on four datasets, resulting in 12 runs. For each run in both 0-shot and 2-shot setups, we compare the model performance averaged over 10 random templates.
Based on the results presented in Table 10, we restricted the final pool of models for evaluation to those that have a consistent increase in performance in the 2-shot setting, in other words, to those demonstrating a performance boost from ICL. More specifically, we kept the models that had 8 or more wins in 2-shot evaluation against 0-shot.
@@ Appendix B Model Selection
Table 10: Evaluation of advanced prediction methods for all models on 4 datasets in 0-shot and 2-shot with random demonstrations. Models that were removed from further evaluation are highlighted in gray. “Calib.” stands for the Calibration prompting method.
@@ Appendix C Full Baseline Results
Table 11 shows the results of evaluation of all 19 models in the default setting with a varying number of few-shot examples. These results illustrate that the template sensitivity issue is present in all models regardless of their size, and is not efficiently mitigated with the increase in the number of demonstrations.
@@
Table 11: Classification accuracy of all models on all datasets in the default setting. The results are aggregated over 10 templates for each set of random demonstrations, i.e., 10 runs for 0-shot and 30 runs for few-shot learning.
@@ Appendix D Template Parts Analysis
(a) Direct prediction method.
(b) Channel prediction method.
(c) Calibration prediction method.
Figure 7: Accuracy for evaluation of templates with fixed parts on the SST-2 dataset with Random 2-shot for all prediction methods.
In addition to studying prompt format sensitivity in general, we analyze how each part of a template impacts model performance. For instance, it could be possible that the inclusion of a certain verbalizer in a template consistently leads to a decline in accuracy, irrespective of the other components.
To find that out, we decompose all templates into their parts and measure the distribution of scores for different variations of each component separately. The results presented in  Figure 7 illustrate that even for state-of-the-art models, such as Llama 2 70B and Falcon 40B, many components exhibit high variance; also, the variance differs between two models. In other words, even if a certain template yields good performance and low variance for a given setup, it is not guaranteed to work consistently well in other setups, and changing a single component could have detrimental effects.
Along with the non-transferability of whole templates, we notice that individual components also do not transfer both between models and prediction methods. For instance, while “It was {}” ranks highest among output verbalizers for Llama 2 70B with the Direct prediction method, it is one of the worst for Falcon 40B.
@@
In summary, there is a complex interaction between the components of a template and their influence on model performance. We hypothesize that the transfer of both whole prompt templates and their parts is limited and requires further analysis.
@@ Appendix E Prediction Methods
We provide the results of advanced prediction methods evaluation for all models in 0-shot and 2-shot setting with random demonstrations in Table 10. We conclude from this comparison that neither of the advanced prediction strategies do not decrease prompt format sensitivity consistently across models and datasets. Moreover, when accounting for the spread in accuracy scores caused by this issue, the advantages of these methods over Direct become less apparent.
Appendix F Reproduction of Results for Advanced Selection Methods
To evaluate the sensitivity of example selection methods to the template choice, we compare how the results reported in original works on these methods change when evaluated on a set of random templates instead of a predefined single one. For an accurate reproduction of original setups, we evaluate both z-ICL and ITM using Channel prompting with corresponding templates from Tables 6 and 7.
In Tables 12 and 13, we present the scores for a single template used in original implementations in the column “Reproduced”, aggregating results over 5 example selection seeds. The “Random” column shows average scores for fixed demonstrations on a set of 10 random templates.
We observe that the results obtained in our code differ from the ones reported in the papers presenting both methods (the “Paper” column). The cause of this discrepancy is presumably the difference in tokenization during preprocessing of the datasets. Both methods use the same codebase with tokenization specific to GPT tokenizers, which results in a significant drop in quality for OPT and LLaMA models. By contrast, our tokenization approach is more general and preserves the ICL performance of these models.
Model SST 2 DBPedia 14
Paper Reproduced Random Paper Reproduced Random
GPT2-Large 0.860.01 0.840.02 0.800.11 0.570.03 0.620.05 0.630.07
GPT2-XL 0.830.04 0.730.13 0.750.12 0.590.03 0.610.02 0.620.09
GPT-J 0.880.02 0.870.02 0.860.06 0.600.04 0.540.05 0.710.05
OPT-6.7B 0.740.03 0.860.03 0.870.04 0.290.02 0.620.04 0.720.06
LLaMA-7B 0.610.05 0.810.12 0.870.07 0.170.01 0.680.02 0.800.06
Table 12: Mean accuracy and standard deviation of 4-shot learning with ITM demonstrations using the Channel prediction method.
Model Paper Reproduced Random
GPT-J 0.830.02 0.860.00 0.610.06
GPT-NeoX 0.790.00 0.830.00 0.800.04
Table 13: Mean accuracy and standard deviation of 4-shot learning with z-ICL demonstrations on SST-2 dataset using the Channel prediction method.
From these results, we conclude that both methods are not robust to the template choice, as the mean performance decreases for multiple models while the standard deviation across seeds increases. Therefore, the gains from advanced example selection methods are caused to a certain degree by the choice of a proper prompt format rather than the retrieved demonstrations.
Appendix G Example Selection Methods
Full results of evaluation of demonstration selection techniques in 4-shot learning using the Direct prediction method are presented in Table 14.
@@ Appendix G Example Selection Methods
The results highlight that advanced example selection techniques often perform comparably to the random choice baseline when evaluated on multiple templates. One might argue that the prompt format choice is inseparable from the method itself and thus such a comparison is invalid. However, since the best-performing formats do not transfer between models or demonstration sets of different sizes selected with the same method, a proper evaluation would require finding the best template for each setup. This procedure both is computationally expensive and difficult to accomplish, as authors of example selection methods frequently omit the description of their format selection algorithm.
Appendix H Accuracy as a Function of Template Rank
We plot the dependence of accuracy on the rank of the template in Figure 8. The results are aggregated across 19 models. Each model was evaluated on 30 random templates with the Direct prediction method and the same set of 2 randomly selected demonstrations. We observe that for SST-2 and AGNews datasets, the mean quality of the tenth-best template is within 0.9 of the best template score, which we consider a successful transfer. Despite the more rapid decay for DBPedia and TREC, taking variation across models into account, we still count first 10 formats as performing on par with the best one.
Figure 8: Relative quality of templates sorted by their classification accuracy. The shaded area indicates the standard deviation across 19 models.
Appendix I Transfer Evaluation with Spearman Rank Correlation
One of the possible means to evaluate template transfer is to calculate the Spearman rank correlation between scores of all templates. As can be seen from Figure 9, this method yields higher correlations than IoU over 10 best formats, but the capacity for transfer is still far from perfect (for example, for SST-2 and TREC datasets).
Figure 9: Spearman rank correlation over 30 templates for models evaluated in the Direct-Random-2-shot setting.
Appendix J IoU Transfer For All Datasets
@@ Appendix I Transfer Evaluation with Spearman Rank Correlation
Similarly to Figure 4, we provide Intersection-over-Union of 10 best prompt formats for all 19 models and all datasets explored in our work in Figure 10. These heatmaps illustrate that the transfer of best-performing templates between models is remarkably low for all datasets.
Appendix K Additional Results For Template Ensembles
Tables 15 and 16 show the results of Template Ensembles evaluation on a broader set of models and datasets. For most setups, ensemble of size 5 exhibit better performance than a single template.
@@
Table 16: Comparison of 2-shot learning performance on the TREC dataset using ensembles of 5 templates and a single template. Results are averaged over 5 random seeds.
Appendix L Evaluation of Instruct Models
We validate that our findings hold true even for the latest instruction-tuned models, such as Llama 3 8B Instruct and Mistral v0.3 7B Instruct. First, as we show in Table 17, in the baseline setting, an increase in number of demonstrations generally leads to a better performance of instruct models but does not significantly decrease variance of their final scores, similarly to what we observe in base models.
Second, Table 18 demonstrates that, after adjusting to template robustness, the default Direct prediction method performs on par with more advanced methods, e.g. Channel and Calibration, while sometimes having a less variance of the model’s scores.
Finally, we observe that the examples retrieved by z-ICL method turn out to be consistently worse than two other methods. This is also in line with our observation for base models. However, in contrast to our main results, where there was no evident winner between two other methods, the demonstrations selected with ITM method turn out to be slightly more robust to template choice for instruction-tuned models, as we show in Table 19.
Model SST-2 DBPedia AGNews TREC
0 2 4 0 2 4 0 2 4 0 2 4
Mistral v0.3 7B Instruct 0.830.07 0.900.09 0.910.12 0.500.06 0.640.05 0.650.02 0.340.05 0.400.08 0.520.08 0.710.07 0.820.07 0.760.08
Llama 3 8B Instruct 0.770.12 0.870.13 0.910.10 0.480.02 0.550.13 0.630.05 0.240.06 0.390.09 0.480.11 0.770.06 0.770.18 0.780.10
Table 17: Classification accuracy of instruction-tuned models on all datasets in the default setting. The results are averaged over 10 templates for each set of random demonstrations (3 sets in total), i.e., 10 runs for 0-shot and 30 runs for few-shot learning.
Model N SST-2 DBPedia AGNews TREC
Mistral v0.3 7B Instruct 0 0.830.07 0.730.05 0.830.10 0.500.06 0.470.08 0.750.09 0.340.05 0.190.08 0.300.07 0.710.07 0.590.06 0.680.06
2 0.900.09 0.830.09 0.900.12 0.640.05 0.750.04 0.910.07 0.400.08 0.410.06 0.430.06 0.820.07 0.800.02 0.800.04
Llama 3 8B Instruct 0 0.770.12 0.660.04 0.740.14 0.480.02 0.480.05 0.720.08 0.240.06 0.180.08 0.220.04 0.770.06 0.570.06 0.760.06
2 0.870.13 0.780.10 0.900.09 0.550.13 0.720.04 0.850.11 0.390.09 0.360.09 0.440.06 0.770.18 0.720.10 0.800.08
Table 18: Evaluation of advanced prediction methods for instruction-tuned models on 4 datasets in 0-shot and 2-shot with random demonstrations. “Calib.” stands for the Calibration prompting method.
Model SST-2 DBPedia AGNews TREC
Random ITM z-ICL Random ITM z-ICL Random ITM z-ICL Random ITM z-ICL
Mistral v0.3 7B Instruct 0.910.12 0.910.10 0.860.08 0.650.02 0.650.06 0.370.05 0.520.08 0.470.10 0.370.02 0.760.08 0.830.06 0.670.06
Llama 3 8B Instruct 0.910.10 0.920.08 0.770.08 0.630.05 0.630.08 0.260.05 0.480.11 0.450.09 0.260.04 0.780.10 0.860.04 0.530.10
Table 19: Evaluation of advanced selection methods for instruction-tuned models using Direct prediction method in 4-shot. Results are aggregated over 10 random templates for each of the three demonstrations selection seeds.
L.1 Ensemble Results
Table 20 shows that applying Template Ensembles method to instruct models results in improved mean classification accuracy with significantly reduced variance across different templates in all configurations that we test.
Model Direct Channel Calibration
Single Ensemble Single Ensemble Single Ensemble
SST-2
Llama 3 8B Instruct 0.870.13 0.940.01 0.780.10 0.860.02 0.900.09 0.950.00
Mistral v0.3 7B Instruct 0.900.09 0.940.01 0.830.09 0.910.02 0.900.12 0.950.00
DBPedia
Llama 3 8B Instruct 0.560.12 0.640.06 0.720.04 0.790.01 0.850.11 0.950.02
Mistral v0.3 7B Instruct 0.640.05 0.670.02 0.750.04 0.830.02 0.910.07 0.950.05
TREC
Llama 3 8B Instruct 0.390.09 0.470.01 0.360.09 0.430.03 0.440.06 0.490.02
Mistral v0.3 7B Instruct 0.400.08 0.480.02 0.410.06 0.460.04 0.430.06 0.470.03
AGNews
Llama 3 8B Instruct 0.770.18 0.890.02 0.720.10 0.810.01 0.800.08 0.840.01
Mistral v0.3 7B Instruct 0.820.07 0.860.02 0.800.02 0.830.01 0.800.04 0.820.02
Table 20: Comparison of 2-shot learning performance of Template Ensembles against other prediction methods on all datasets used in our work. We use ensembles of size 5 and average the results over 5 random seeds